what percent of 55 is 22
Answer by josmiceli(19441)

" per cent " means " per one-hundred "
( in latin ), so
" what / 100 = 22/55 "
X/100 = 22/55
X = ( 22/55 ) * 100
X = 20
22 is 20 % of 55